# Two-Step Data Augmentation for Masked Face Detection and Recognition: Turning Fake Masks to Real

### v1

@@ 1 INTRODUCTION
Computer Vision tasks like recognition, detection, classification, etc., performed on occluded human faces, existed even before the COVID-19 outbreak (8099536). The spread of COVID-19 has urgently imposed performance and robustness requirements to such applications. However, multiple latest masked-face detection or recognition works (singh2021face; DBLP:journals/corr/abs-2104-09874) claim their models only as starting points of future transfer learning with more data, instead of final results, or at least recognize the data insufficiency problem. The research community has built mature datasets of dominantly full faces, but masked face datasets are still under construction.
For full faces, recognition tasks use single-face image sets, with every single identity assigned to multiple facial images (8614364), and detection algorithms work on multi-face scene images and learn bounding box locations together with an optional class label for each box (fddbTech; 8569109). To improve the masked-face datasets to facilitate the same learning approaches, we either collect and annotate raw data or generate artificial images to augment existing data (2020). In this paper, we focus on generating artificially masked faces.
Recent works address this problem by modifying unmasked faces into masked. Some use rules to warp masks onto faces (wang2020masked; DBLP:journals/corr/abs-2008-11104; cabani2021maskedface). One uses Neural Network (NN) to translate unmasked faces into masked (geng2020masked). The rule-based methods provide realistic mask textures and completely avoid the risk of distorting other parts of faces. However, they often result in bad transitions between masks and background faces, and the lighting on masks often looks unreal. Some rule-based algorithms (DBLP:journals/corr/abs-2008-11104; DBLP:journals/corr/abs-2109-05804) achieved mask diversity by defining different mask image options, but this diversity is limited to the predefined mask types. In contrast, NN methods learn to avoid facial distortions and generate mask textures, often to a reasonable extent but never with promise. They provide natural transition, realistic details, and sometimes more diversity in compensation.
@@ 2 RELATED WORK
Real-World Datasets. Traditional full-face datasets embody a large number of faces with high variations in demographics, head rotations, facial landmarks, occlusion degrees, facial expressions, etc. (SAGONAS20163; DBLP:journals/corr/abs-1812-04948)  [^1] Occlusions here include various types like body parts, scarves, etc. Medical masks constitute only a tiny portion. Annotations to these datasets include categories mentioned above, as well as subject identities for facial recognition, bounding boxes for facial detection, and so on.
@@
On the other hand, datasets emphasizing masked faces are limited in the quantity of data and the variation of features. wang2020masked developed two different datasets based on real-world images. Real-World Masked Face Dataset (RMFD) (wang2020masked) for recognition tasks contains 5,000 masked and 90,000 normal faces belonging to 525 people. Masked Face Detection Dataset (MFDD) (wang2020masked) contains 24,711 masked face images for detection tasks, which is currently not publicly accessible  [^2] https://github.com/X-zhangyang/Real-World-Masked-Face-Dataset/issues/16. MAsked FAces (MAFA) by 8099536 with 30,811 internet images and 35,806 masked faces is the largest real-world masked face dataset to our best knowledge. It is annotated with relatively abundant information such as face bounding boxes, mask bounding boxes, mask types, face orientations, occlusion degrees, gender, race, and more, but MAFA alone is not comparable with all the diversity of currently available full-face datasets. Moreover, no identity information is provided to MAFA, so it is only for face detection tasks. Besides RMFD, recognition tasks may also use Masked Faces in Real World for Face Recognition (MFR2) (DBLP:journals/corr/abs-2008-11104), a small set of 269 images belonging to 53 identities. geng2020masked proposed a two-part Masked Face Segmentation and Recognition (MFSR) dataset. 9,742 masked images were collected from the web and annotated with mask regions in the first part. The second part contains 11,615 faces, masked or non-masked, for 704 real-world identities and 300 internet-obtained identities.
@@
Artificially Masked Faces. Some researchers generated artificial data to augment existing real-world masked face datasets. Stemmed from RMFD and MFDD, the same lab proposed Simulated Masked Face Recognition Dataset (SMFRD) (wang2020masked), using a naive copy-and-paste method to put cartoon mask images onto existing face recognition datasets. DBLP:journals/corr/abs-2008-11104 and cabani2021maskedface separately used more sophisticated methods to warp mask images onto faces based on detected facial landmarks (SAGONAS20163). DBLP:journals/corr/abs-2008-11104 provided their method as a MaskTheFace tool for both single- and multi-face images. cabani2021maskedface published a single-face dataset, MaskedFace-Net, applying their rule-based method to Flickr-Faces-HQ (DBLP:journals/corr/abs-1812-04948). MaskedFace-Net consists of Correctly and Incorrectly Masked Face Dataset (CMFD and IMFD) with about 70,000 synthesized images each. CMFD is used as our model input. While our paper was in progress, DBLP:journals/corr/abs-2109-05804 published MLFW (Masked LFW), which enhanced the landmark-based warping by extra rule-based improvements on the unreal lighting and abrupt mask boundaries.
@@
Beyond rule-based methods, geng2020masked proposed an NN model, Identity Aware Mask GAN (IAMGAN), to synthesize masked faces. It consists of a CycleGAN-like generation module and an Identity Preservation (IP) module. The IP module has a mask region predictor and an identity classifier. The former predicts mask regions and removes them from both masked (output) and unmasked (input) images. The latter classifies identities using the rest parts and penalizes identity change between input and output based on information from multiple classifier levels.
@@
GAN Models for I2I. Image to image translation is a problem of mapping images from one domain into another, with the intrinsic source content preserved and the extrinsic target style transferred (DBLP:journals/corr/abs-2101-08629). It has been used in a broad set of scenarios such as image synthesis (DBLP:journals/corr/abs-1803-03396), image segmentation (GUO2020127), style transfer (DBLP:journals/corr/ZhuPIE17), and more. GAN models are adapted to the I2I tasks by replacing the generator’s standard random input with source image embeddings and keeping the discriminator’s function of distinguishing synthesized and authentic images.
@@
The earliest versions of GANs for image translation, such as pix2pix (DBLP:journals/corr/IsolaZZE16), require aligned image pairs as the training data. However, getting a large number of strictly aligned image pairs is hard or even impossible in many situations. DBLP:journals/corr/ZhuPIE17 proposed CycleGAN using a cyclic loss, making it possible to train an I2I on unpaired data. The previously mentioned IAMGAN (geng2020masked) is an enhanced CycleGAN aiming specifically to generate masks. Based similarly on CycleGAN, tang2021attentiongan trained the generator to produce attention masks together with the generated image contents, which guided the generator to translate individual objects without changing the background. They call their general-purpose model “AttentionGAN”.
@@
Our work adapts AttentionGAN (tang2021attentiongan). The key innovation is that our source images are not full-face but rule-based mask images, i.e., CMFD (cabani2021maskedface). We use a rule-based method as the first step and the adapted AttentionGAN as the second step. We show that: 1. warping mask images onto full faces provides style guidance and ground truth attention for better GAN model results; 2. the GAN model renders the rule-based results into more realistic details. With the latest advances in rule-based methods such as DBLP:journals/corr/abs-2109-05804, our two-step proposal may still improve details.
@@ 4.2 Adding Noise
Inspired by StyleGAN (DBLP:journals/corr/abs-1812-04948), which takes random noise input to multiple generator layers for result diversity, we modified AttentionGAN to include zero-mean Gaussian noise input of a similar style. We first attempted noise input to the first content-generating TC layer following the nine ResNet blocks. We tested the noise-tuned model structure on parameters trained without noise and found that the generated images remained the same no matter how we amplified the noise. We concluded that the first TC layer assimilates noise and does not map noise input to identifiable features in the final output. Then, with one TC layer noise-tuned at a time, we tested different amplitudes of noise input on the other two content-generating TC layers, using the same model mentioned above. We found that noise added to either of the other two TC layers makes a difference to the model output, producing random noise pixels as shown in Figure 3.
@@ 4.3 Transfer Learning and the Training Timeline
Our initial parameters for the single-face mask generation task are from a different research topic: a multi-face image translator trained with the same AttentionGAN model. These initial weights were gained by 60 epochs of training on Face Detection Data Set and Benchmark (fddbTech) and MAFA data, with hyperparameters in the original AttentionGAN code set to $lambda_{A}=5$, $lambda_{B}=5$, and $lambda_{identity}=0.2$. From each of FDDB and MAFA, about 25,000 images were selected as training sets A and B separately. We will not elaborate on the rationale behind our choices since it belongs to other research. The information here is only for the reproduction of our result.
@@ 5.2 Comparison with IAMGAN
Both geng2020masked and our research used CycleGAN-based methods to turn full faces into masked ones. Our differences include:
@@
Performance scores such as Frechet Inception Distance (FID) (DBLP:journals/corr/HeuselRUNKH17) and Kernel Inception Distance (KID) (binkowski2018demystifying) are usually used to compare the fidelity of different synthesized datasets, but we lack a real-world masked face dataset as the baseline. Datasets highly similar to either IAMGAN’s or our training data would unfaithfully push one party’s score high. Therefore, we ran our model on some examples shown in the IAMGAN paper (geng2020masked) and demonstrated a qualitative comparison in Figure 9.
@@ 5.3 Potential Improvements
Future work on datasets may also emphasize increasing mask color/type diversity, but the type diversity should be attempted together with increased model abilities in learning diverse mask shapes. The cycle loss in our model is better at dealing with point-to-point mapping with little shape-changing, so it may be insufficient for the shape diversity. Besides, we only need the A-to-B model, not the opposite direction. Therefore, the single-sided domain mapping proposed by DBLP:journals/corr/BenaimW17, with its distance constraints substituting the cycle loss, could be one direction we consider together with an increased mask type diversity. If we retain the two-directional training architecture, the distance constraints may even be used together with the cycle loss.
@@ 6 CONCLUSIONS
We aimed at turning full face detection/recognition datasets into masked face datasets, supplementing the limited training data for masked face tasks. For this purpose, we proposed a two-step data augmentation method, utilizing cabani2021maskedface’s algorithm to warp mask images onto faces as a pre-step to an AttentionGAN-like model that generates more realistically masked faces. We applied multiple improvements to the GAN model training and verified their effectiveness through experimental results. Analyses of our final results showed that the two-step method provided noticeable improvements compared to using a rule-based method alone. Even with the latest advances of the rule-based method by DBLP:journals/corr/abs-2109-05804, we still expect an extra I2I step to render the rule-based results with more details, such as irregular region boundaries caused by fabric folds and straps. Our results are also comparable with state-of-the-art NN-only mask generation methods such as IAMGAN, with complementary details. For example, we produced lighting changes and mask stripes or their connecting points missing in IAMGAN results.

### v2

@@ 1 INTRODUCTION
Computer Vision tasks like recognition, detection, classification, etc., performed on occluded human faces, existed even before the COVID-19 outbreak (Ge et al., 2017). The spread of COVID-19 has urgently imposed performance and robustness requirements to such applications. However, multiple latest masked-face detection or recognition works (Singh et al., 2021; Montero et al., 2021) claim their models only as starting points of future transfer learning with more data, instead of final results, or at least recognize the data insufficiency problem. The research community has built mature datasets of dominantly full faces, but masked face datasets are still under construction.
For full faces, recognition tasks use single-face image sets, with every single identity assigned to multiple facial images (Masi et al., 2018), and detection algorithms work on multi-face scene images and learn bounding box locations together with an optional class label for each box (Jain and Learned-Miller, 2010; Yang and Jiachun, 2018). To improve the masked-face datasets to facilitate the same learning approaches, we either collect and annotate raw data or generate artificial images to augment existing data (Wang et al., 2020a). In this paper, we focus on generating artificially masked faces.
Recent works address this problem by modifying unmasked faces into masked. Some use rules to warp masks onto faces (Wang et al., 2020b; Anwar and Raychowdhury, 2020; Cabani et al., 2021). One uses Neural Network (NN) to translate unmasked faces into masked (Geng et al., 2020). The rule-based methods provide realistic mask textures and completely avoid the risk of distorting other parts of faces. However, they often result in bad transitions between masks and background faces, and the lighting on masks often looks unreal. Some rule-based algorithms (Anwar and Raychowdhury, 2020; Wang et al., 2021) achieved mask diversity by defining different mask image options, but this diversity is limited to the predefined mask types. In contrast, NN methods learn to avoid facial distortions and generate mask textures, often to a reasonable extent but never with promise. They provide natural transition, realistic details, and sometimes more diversity in compensation.
@@ 2 RELATED WORK
Real-World Datasets. Traditional full-face datasets embody a large number of faces with high variations in demographics, head rotations, facial landmarks, occlusion degrees, facial expressions, etc. (Sagonas et al., 2016; Karras et al., 2018)  [^1] Occlusions here include various types like body parts, scarves, etc. Medical masks constitute only a tiny portion. Annotations to these datasets include categories mentioned above, as well as subject identities for facial recognition, bounding boxes for facial detection, and so on.
@@
On the other hand, datasets emphasizing masked faces are limited in the quantity of data and the variation of features. Wang et al. (2020b) developed two different datasets based on real-world images. Real-World Masked Face Dataset (RMFD) (Wang et al., 2020b) for recognition tasks contains 5,000 masked and 90,000 normal faces belonging to 525 people. Masked Face Detection Dataset (MFDD) (Wang et al., 2020b) contains 24,711 masked face images for detection tasks, which is currently not publicly accessible  [^2] https://github.com/X-zhangyang/Real-World-Masked-Face-Dataset/issues/16. MAsked FAces (MAFA) by Ge et al. (2017) with 30,811 internet images and 35,806 masked faces is the largest real-world masked face dataset to our best knowledge. It is annotated with relatively abundant information such as face bounding boxes, mask bounding boxes, mask types, face orientations, occlusion degrees, gender, race, and more, but MAFA alone is not comparable with all the diversity of currently available full-face datasets. Moreover, no identity information is provided to MAFA, so it is only for face detection tasks. Besides RMFD, recognition tasks may also use Masked Faces in Real World for Face Recognition (MFR2) (Anwar and Raychowdhury, 2020), a small set of 269 images belonging to 53 identities. Geng et al. (2020) proposed a two-part Masked Face Segmentation and Recognition (MFSR) dataset. 9,742 masked images were collected from the web and annotated with mask regions in the first part. The second part contains 11,615 faces, masked or non-masked, for 704 real-world identities and 300 internet-obtained identities.
@@
Artificially Masked Faces. Some researchers generated artificial data to augment existing real-world masked face datasets. Stemmed from RMFD and MFDD, the same lab proposed Simulated Masked Face Recognition Dataset (SMFRD) (Wang et al., 2020b), using a naive copy-and-paste method to put cartoon mask images onto existing face recognition datasets. Anwar and Raychowdhury (2020) and Cabani et al. (2021) separately used more sophisticated methods to warp mask images onto faces based on detected facial landmarks (Sagonas et al., 2016). Anwar and Raychowdhury (2020) provided their method as a MaskTheFace tool for both single- and multi-face images. Cabani et al. (2021) published a single-face dataset, MaskedFace-Net, applying their rule-based method to Flickr-Faces-HQ (Karras et al., 2018). MaskedFace-Net consists of Correctly and Incorrectly Masked Face Dataset (CMFD and IMFD) with about 70,000 synthesized images each. CMFD is used as our model input. While our paper was in progress, Wang et al. (2021) published MLFW (Masked LFW), which enhanced the landmark-based warping by extra rule-based improvements on the unreal lighting and abrupt mask boundaries.
@@
Beyond rule-based methods, Geng et al. (2020) proposed an NN model, Identity Aware Mask GAN (IAMGAN), to synthesize masked faces. It consists of a CycleGAN-like generation module and an Identity Preservation (IP) module. The IP module has a mask region predictor and an identity classifier. The former predicts mask regions and removes them from both masked (output) and unmasked (input) images. The latter classifies identities using the rest parts and penalizes identity change between input and output based on information from multiple classifier levels.
@@
GAN Models for I2I. Image to image translation is a problem of mapping images from one domain into another, with the intrinsic source content preserved and the extrinsic target style transferred (Pang et al., 2021). It has been used in a broad set of scenarios such as image synthesis (Regmi and Borji, 2018), image segmentation (Guo et al., 2020), style transfer (Zhu et al., 2017), and more. GAN models are adapted to the I2I tasks by replacing the generator’s standard random input with source image embeddings and keeping the discriminator’s function of distinguishing synthesized and authentic images.
@@
The earliest versions of GANs for image translation, such as pix2pix (Isola et al., 2016), require aligned image pairs as the training data. However, getting a large number of strictly aligned image pairs is hard or even impossible in many situations. Zhu et al. (2017) proposed CycleGAN using a cyclic loss, making it possible to train an I2I on unpaired data. The previously mentioned IAMGAN (Geng et al., 2020) is an enhanced CycleGAN aiming specifically to generate masks. Based similarly on CycleGAN, Tang et al. (2021) trained the generator to produce attention masks together with the generated image contents, which guided the generator to translate individual objects without changing the background. They call their general-purpose model “AttentionGAN”.
@@
Our work adapts AttentionGAN (Tang et al., 2021). The key innovation is that our source images are not full-face but rule-based mask images, i.e., CMFD (Cabani et al., 2021). We use a rule-based method as the first step and the adapted AttentionGAN as the second step. We show that: 1. warping mask images onto full faces provides style guidance and ground truth attention for better GAN model results; 2. the GAN model renders the rule-based results into more realistic details. With the latest advances in rule-based methods such as Wang et al. (2021), our two-step proposal may still improve details.
@@ 4.2 Adding Noise
Inspired by StyleGAN (Karras et al., 2018), which takes random noise input to multiple generator layers for result diversity, we modified AttentionGAN to include zero-mean Gaussian noise input of a similar style. We first attempted noise input to the first content-generating TC layer following the nine ResNet blocks. We tested the noise-tuned model structure on parameters trained without noise and found that the generated images remained the same no matter how we amplified the noise. We concluded that the first TC layer assimilates noise and does not map noise input to identifiable features in the final output. Then, with one TC layer noise-tuned at a time, we tested different amplitudes of noise input on the other two content-generating TC layers, using the same model mentioned above. We found that noise added to either of the other two TC layers makes a difference to the model output, producing random noise pixels as shown in Figure 3.
@@ 4.3 Transfer Learning and the Training Timeline
Our initial parameters for the single-face mask generation task are from a different research topic: a multi-face image translator trained with the same AttentionGAN model. These initial weights were gained by 60 epochs of training on Face Detection Data Set and Benchmark (Jain and Learned-Miller, 2010) and MAFA data, with hyperparameters in the original AttentionGAN code set to $lambda_{A}=5$, $lambda_{B}=5$, and $lambda_{identity}=0.2$. From each of FDDB and MAFA, about 25,000 images were selected as training sets A and B separately. We will not elaborate on the rationale behind our choices since it belongs to other research. The information here is only for the reproduction of our result.
@@ 5.2 Comparison with IAMGAN
Both Geng et al. (2020) and our research used CycleGAN-based methods to turn full faces into masked ones. Our differences include:
@@
Performance scores such as Frechet Inception Distance (FID) (Heusel et al., 2017) and Kernel Inception Distance (KID) (Bińkowski et al., 2018) are usually used to compare the fidelity of different synthesized datasets, but we lack a real-world masked face dataset as the baseline. Datasets highly similar to either IAMGAN’s or our training data would unfaithfully push one party’s score high. Therefore, we ran our model on some examples shown in the IAMGAN paper (Geng et al., 2020) and demonstrated a qualitative comparison in Figure 9.
@@ 5.3 Potential Improvements
Future work on datasets may also emphasize increasing mask color/type diversity, but the type diversity should be attempted together with increased model abilities in learning diverse mask shapes. The cycle loss in our model is better at dealing with point-to-point mapping with little shape-changing, so it may be insufficient for the shape diversity. Besides, we only need the A-to-B model, not the opposite direction. Therefore, the single-sided domain mapping proposed by Benaim and Wolf (2017), with its distance constraints substituting the cycle loss, could be one direction we consider together with an increased mask type diversity. If we retain the two-directional training architecture, the distance constraints may even be used together with the cycle loss.
@@ 6 CONCLUSIONS
We aimed at turning full face detection/recognition datasets into masked face datasets, supplementing the limited training data for masked face tasks. For this purpose, we proposed a two-step data augmentation method, utilizing Cabani et al. (2021)’s algorithm to warp mask images onto faces as a pre-step to an AttentionGAN-like model that generates more realistically masked faces. We applied multiple improvements to the GAN model training and verified their effectiveness through experimental results. Analyses of our final results showed that the two-step method provided noticeable improvements compared to using a rule-based method alone. Even with the latest advances of the rule-based method by Wang et al. (2021), we still expect an extra I2I step to render the rule-based results with more details, such as irregular region boundaries caused by fabric folds and straps. Our results are also comparable with state-of-the-art NN-only mask generation methods such as IAMGAN, with complementary details. For example, we produced lighting changes and mask stripes or their connecting points missing in IAMGAN results.